OpenStudy (anonymous):

Jerome is going to draw a marble from the bag shown below, replace it, and then draw another marble. What is the probability that Jerome will get a green or yellow marble on the first draw and a pink on the second draw?

OpenStudy (anonymous):

(3 green, 4 yellow, 5 blue, 8 pink)

OpenStudy (anonymous):

7/50 3/8 27/200

OpenStudy (kropot72):

Are the three probabilities you have posted choices of answer?

OpenStudy (anonymous):

answers

OpenStudy (kropot72):

Probability of drawing a green marble is 3/20 Probab ility of drawing a yellow marble is 4/20 What do you think will be the probability of drawing either a green or a yellow marble?

OpenStudy (anonymous):

7/20

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Correct. Good thinking :) Now what is the probability of drawing a pink marble? Don't forget that all 20 marbles are in the bag.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8/20

OpenStudy (kropot72):

You are soo right. Good work :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do i do next\

OpenStudy (kropot72):

You have both answers. Strange that the choices you gave earlier don't include them. Can you check your postings of the question and the choices of answer.

OpenStudy (anonymous):

thats all of them

OpenStudy (kropot72):

Well assuming your posting of the question is correct, none of the choices is correct. I am sure of the answers 7/20 and 8/20 = 2/5.

OpenStudy (anonymous):

My result is 7/50 !

OpenStudy (anonymous):

I think she miscalculated somewhere!

OpenStudy (anonymous):

how

OpenStudy (anonymous):

( 3/20 * 8/20 ) + ( 1/5 * 8/20 )

OpenStudy (anonymous):

oh ok thank you

OpenStudy (kropot72):

Sorry, I took them as separate probabilities which are correct as I gave them. the combined probability is (7/20) * (8/20) = 7/50